# What Makes a Language Easy to Deep-Learn?

### v3

@@ 4 Methods
Crucially, the ten input languages have different degrees of structure, ranging from languages with no structure to languages with consistent, systematic grammar. Each language has a structure score represented by topographic similarity [15], quantifying the degree to which similar labels describe similar meanings. The topographic similarity is measured as the correlation between all labels’ pairwise length-normalized edit distances and their corresponding pairwise semantic differences. The semantic difference between two scenes is calculated as the sum of the difference in shape and the difference in angles [27]. The difference in shape is zero if the two scenes contain the same shape, and one otherwise. The difference in angles is calculated as the absolute difference divided by 180. The topographic similarity of a language is then calculated as the pairwise correlation between all semantic differences and all normalized edit distances. For a complete list of input languages and their structure scores, see Appendix A.
@@
For the large language models, we supplied the full training data of the respective input language to GPT-3: 23 lines consisting of shape-angle pairs in a textual format, and the corresponding target label. These 23 lines were followed by a single line that only contained shape and angle but no word. GPT-3 was made to predict the most likely word as completion, for which it could take into account the 23 triples presented in the prompt. In the memorization task, the target word appears earlier in the prompt, which means that the perfect solution would be to simply copy this word. In the generalization task, we gave GPT-3 a combination of shape and angle not present in the training data (and not in the prompt). The model generated the most likely descriptive word for the new shape-angle pair.
We had to make certain technical choices when using GPT-3. First, we chose a consistent input representation (Javascript Object Notation). We do not insert a task description to avoid potential bias. Instead we purely rely on next-token prediction. Second, we set the sampling temperature to zero, which controls the randomness of the generation, such that we obtain deterministic generations. Third, we do not impose any restrictions on the characters that can be generated but rely on its ability to detect this pattern from the training data. Fourth, we do not feed back GPT-3’s previous predictions into the prompt. Lastly, GPT-3’s tokenization procedure (how text is split into subword tokens) could have been problematic for applying it to our artificial languages. However, we found that GPT-3 still reaches high memorization performance, which suggests that tokenization is not a problem. We have confirmed that the words of the artificial languages are tokenized as expected with the OpenAI’s Tokenizer [^1] https://platform.openai.com/tokenizer: falling back to one token per character.
@@
The output dimension of Encoder, the hidden state sizes of Reader and Writer, and the embedding size are all set to 50. A sensitivity analysis of the hidden size on the dependent variables of interest is provided in Appendix D. Similarly to Nakkiran et al. [57], larger hidden sizes led to a faster increase in memorization and generalization.
@@
In the exposure block, human participants were exposed to scenes with the corresponding target labels. Therefore, we train the deep learning models using a loss function with two terms: a generative and a contrastive loss term. The generative loss, $\mathcal{L}_{\mathrm{gen}}$, is a token-wise cross-entropy with the ground-truth label of the original language. The contrastive loss, $\mathcal{L}_{\mathrm{con}}$, promotes similar latent representations of scenes and labels that correspond to each other and contrasts representations that do not. Specifically, we use the normalized temperature scaled cross-entropy loss (NTXent) [70]. We use other scenes in the same batch as distractors for the contrastive loss term. The final loss function is $\mathcal{L}=\mathcal{L}_{\mathrm{gen}}+\alpha_{\mathrm{con}}\mathcal{L}_{\mathrm{con}}$. The factor $\alpha_{\mathrm{con}}$ determines the relative weight of the loss terms. For the main experiment, we use $\alpha_{\mathrm{con}}=0.1$. A sensitivity analysis using other values for $\alpha_{\mathrm{con}}$ is provided in Appendix D.
In the guessing block, we use the same loss function as in the exposure block. The contrastive loss term $\mathcal{L}_{\mathrm{con}}$ mirrors the task in which human participants had to select the correct scene against the distractors given a label. The generative loss term $\mathcal{L}_{\mathrm{gen}}$ is used so that the model does not “forget” how to generate [93].
In more detail, the latent representation $\bm{z}=\textsc{Encoder}(\bm{x})$ of the scene $\bm{x}$ should be closest to the latent representation $\bm{z}^{\prime}=\textsc{Reader}(\bm{m})$ of the corresponding label $\bm{m}$. The difference from exposure training is that in the guessing block, we use the identical distractors used in experiments with humans, whereas, in the exposure block, we use the other scenes from the same batch. The trajectory of guessing accuracy during training is shown in Appendix B.
@@
The parameters are randomly initialized by He initialization [94], the default initialization method in PyTorch [95]. We employ the widely used Adam optimizer [96] to carry out the optimization of the loss function with the default learning rate of $10^{-3}$. As common in machine learning, we have to make certain decisions about the neural network architecture design, optimization procedure, and hyperparameters. All these decisions may impact the results. However, we have varied relevant hyperparameter settings and found that the results are robust and do not dependent on specific settings of the hyperparameters (see Appendix D).
@@
Production similarity measures the overlap between two sets of labels. It is computed as one minus the normalized edit distance between pairs of labels. For our analysis, we use production similarity once to quantify the similarity between the generated labels and the ground truth of the input languages, and once to quantify the similarity of labels generated by neural network agents with labels produced by human learners.
The generalization score measures the degree of systematicity during the generalization test [27]. We take two sets of scenes: a training set, on which the agents were trained, and a test set, on which the agents were not trained. We then do the following for each agent. First, we take two sets of labels: one previously generated for each training scene by the agent and another that we let the agent generate for each test scene. Second, the difference between train and test scenes is measured by pairwise semantic difference. Semantic difference is calculated as described above under Input Languages. Third, the difference between train and test labels is measured by pairwise normalized edit distance. Finally, we compute the Pearson correlation between these two differences across all scenes. Then, we take the average correlation coefficient across all agents as the generalization score.
@@
We trained 100 differently-initialized neural network models over 100 rounds for each of the ten input languages. The testing in each round consisted of 23 memorization and 13 generalization examples. This makes a total of 2.3M memorization and 1.3M generalization test results subject to statistical analyses. Significance was tested using linear mixed-effects models, as implemented in the Python package statsmodels [97], for production similarity (LME 1), generalization score (LME 3), generalization convergence (LME 4), as well as production similarity to humans in memorization (LME 2) and generalization (LME 5). We use the structure score and the logarithmized round number in all measures as a fixed effect. The number of rounds was logarithmized following scaling laws of neural language models [58]. Both the structure score and the logarithmized round number were centered and scaled We consider two random effects: the random seed for initialization (which also determines the input language) and the specific scene. For LME 5, scaling the log-transformed round number to unit variance hindered convergence, so the log rounds were only centered. The full results of the statistical models are provided in Appendix C.
@@ Acknowledgments
We thank Yosef Prat, Tal Simon, Mitja Nikolaus, Marieke Woensdregt, and Adam Kohan for their comments and discussions. We thank Shinje Wu and Marianne de Heer Kloots for sharing their data.
Data Availability Statement
The code for reproducing our experiments is available at https://github.com/lgalke/easy2deeplearn.
Appendix A Details of the input languages
An overview of the structure score and the resulting structure bin of the ten input languages is provided in Table 2. We further report an ambiguity score for each language, calculated as one minus to the ratio between distinct forms and distinct meanings within the language.
@@ Appendix A Details of the input languages
Appendix B Extended Results
Here, we provide the extended results of our experiments with recurrent neural networks trained from scratch including all 100 epochs and additional results for binary accuracy and guessing accuracy. Figure 6 shows the production similarity to ground truth during memorization. Figure 7 shows the production similarity to humans during memorization. Figure 8 shows the systematicity during generalization. Figure 9 shows the convergence score during generalization. Figure 10 shows the production similarity to humans during generalization. Figure 11 shows the binary accuracy with respect to ground truth during memorization. Figure 12 shows the guessing accuracy during training.
@@ Appendix B Extended Results
Appendix C Details of the Statistical Analyses of the Results
C.1 LME 1: Production Similarity during Memorization
C.2 LME 2: Production Similarity to Humans during Memorization
C.3 LME 3: Systematicity during Generalization
C.4 LME 4: Convergence Score during Generalization
C.5 LME 5: Production Similarity to Humans during Generalization
C.6 LME 6: Production Similarity to Ground Truth at Specific Rounds
Table 3: Mixed Linear Model Regression Results. LME 1: Production similarity to ground truth during memorization. LME 2: Production similarity to Humans during Generalization. LME 3: Generalization Score
@@ C.6 LME 6: Production Similarity to Ground Truth at Specific Rounds
Table 4: Mixed Linear Model Regression Results: Production Similarity to Humans during Memorization at a Fixed Round Number
@@ Appendix D Sensitivity to Hyperparameters
D.1 Sensitivity to Hidden Layer Size
@@
D.2 Sensitivity to the Scaling Factor for the Contrastive Loss Term
@@
D.3 Robustness of the structure effect with respect to hyperparameter choices

### v4

@@ Methods
Crucially, the ten input languages have different degrees of structure, ranging from languages with no structure to languages with consistent, systematic grammar. Each language has a structure score represented by topographic similarity [15], quantifying the degree to which similar labels describe similar meanings. The topographic similarity is measured as the Pearson correlation between all labels’ pairwise length-normalized edit distances and their corresponding pairwise semantic differences. The semantic difference between two scenes is calculated as the sum of the difference in shape and the difference in angles [28]. The difference in shape is zero if the two scenes contain the same shape, and one otherwise. The difference in angles is calculated as the absolute difference divided by 180. The topographic similarity of a language is then calculated as the pairwise correlation between all semantic differences and all normalized edit distances. For a complete list of input languages and their structure scores, see Table 3 in the SI.
@@
For the large language models, we supplied the full training data of the respective input language to GPT-3.5: 23 lines consisting of shape-angle pairs in a textual format, and the corresponding target label. These 23 lines were followed by a single line that only contained shape and angle but no word. GPT-3.5 was made to predict the most likely word as completion, for which it could take into account the 23 triples presented in the prompt. In the memorization task, the target word appears earlier in the prompt, which means that the perfect solution would be to simply copy this word. In the generalization task, we gave GPT-3.5 a combination of shape and angle not present in the training data (and not in the prompt). The model generated the most likely descriptive word for the new shape-angle pair.
We had to make certain technical choices when using GPT-3.5. First, we chose a consistent input representation (Javascript Object Notation). We do not insert a task description to avoid potential bias. Instead we purely rely on next-token prediction. Second, we set the sampling temperature to zero, which controls the randomness of the generation, such that we obtain deterministic generations. Third, we do not impose any restrictions on the characters that can be generated but rely on its ability to detect this pattern from the training data. Fourth, we do not feed back GPT-3.5’s previous productions into the prompt. Lastly, GPT-3.5’s tokenization procedure (how text is split into subword tokens) could have been problematic for applying it to our artificial languages. However, we found that GPT-3.5 still reaches high memorization performance, which suggests that tokenization is not a problem. We have confirmed that the words of the artificial languages are tokenized as expected with the OpenAI’s Tokenizer (https://platform.openai.com/tokenizer): falling back to one token per character.
@@
The output dimension of Encoder, the hidden state sizes of Reader and Writer, and the embedding size are all set to 50. A sensitivity analysis of the hidden size on the dependent variables of interest is provided in Figs. 18–20 in the SI. Similarly to Nakkiran et al. [59], larger hidden sizes led to a faster increase in memorization and generalization.
@@
In the exposure block, human participants were exposed to scenes with the corresponding target labels. Therefore, we train the deep learning models using a loss function with two terms: a generative and a contrastive loss term. The generative loss, $\mathcal{L}_{\mathrm{gen}}$, is a token-wise cross-entropy with the ground-truth label of the original language. The contrastive loss, $\mathcal{L}_{\mathrm{con}}$, promotes similar latent representations of scenes and labels that correspond to each other and contrasts representations that do not. Specifically, we use the normalized temperature-scaled cross-entropy loss (NTXent) [73]. We use other scenes in the same batch as distractors for the contrastive loss term. The final loss function is $\mathcal{L}=\mathcal{L}_{\mathrm{gen}}+\alpha_{\mathrm{con}}\mathcal{L}_{\mathrm{con}}$. The factor $\alpha_{\mathrm{con}}$ determines the relative weight of the loss terms. For the main experiment, we use $\alpha_{\mathrm{con}}=0.1$. A sensitivity analysis using other values for $\alpha_{\mathrm{con}}$ is provided in Figs. 21-23 of the SI.
In the guessing block, we use the same loss function as in the exposure block. The contrastive loss term $\mathcal{L}_{\mathrm{con}}$ mirrors the task in which human participants had to select the correct scene against the distractors given a label. The generative loss term $\mathcal{L}_{\mathrm{gen}}$ is used so that the model does not “forget” how to generate [96]. Notably, the guessing task itself could be also carried out by having the models generate a descriptive label for each scene and then select the closest one to the given label in terms of edit distance. However, we opted for optimizing shared parameters through a contrastive loss to ensure that the guessing task would also have an effect on the production task (and vice-versa).
In more detail, the latent representation $\bm{z}=\textsc{Encoder}(\bm{x})$ of the scene $\bm{x}$ should be closest to the latent representation $\bm{z}^{\prime}=\textsc{Reader}(\bm{m})$ of the corresponding label $\bm{m}$. The difference from exposure training is that in the guessing block, we use the identical distractors used in experiments with humans, whereas, in the exposure block, we use the other scenes from the same batch. The trajectory of guessing accuracy during training is shown in Fig. 12 in the SI.
@@
The parameters are randomly initialized by He initialization [97], the default initialization method in PyTorch [98]. We employ the widely used Adam optimizer [99] to carry out the optimization of the loss function with the default learning rate of $10^{-3}$. As common in machine learning, we have to make certain decisions about the neural network architecture design, optimization procedure, and hyperparameters. All these decisions may impact the results. However, we have varied relevant hyperparameter settings and found that the results are robust and do not dependent on specific settings of the hyperparameters (see Figs. 24 and 25 in the SI).
@@
Production similarity measures the overlap between two sets of labels. It is computed as one minus the normalized edit distance between pairs of labels. For our analysis, we use production similarity once to quantify the similarity between the generated labels and the ground truth of the input languages, and once to quantify the similarity of labels generated by neural network agents with labels produced by human learners. For example, a recurrent neural network that produced ’muif-a’ for shape 3 moving in direction 360 degrees would have a high production similarity to the majority of human participants who produced ’muif-i’.
The generalization score measures the degree of systematicity during the generalization test [28]. We take two sets of scenes: a training set, on which the agents were trained, and a test set, on which the agents were not trained. We then do the following for each agent. First, we take two sets of labels: one previously generated for each training scene by the agent and another that we let the agent generate for each test scene. Second, the difference between train and test scenes is measured by pairwise semantic difference. Semantic difference is calculated as topographic similarity. Third, the difference between train and test labels is measured by pairwise normalized edit distance. Finally, we compute the Pearson correlation between these two differences across all scenes. Then, we take the average correlation coefficient across all agents as the generalization score.
@@
We trained 100 differently-initialized neural network models over 100 rounds for each of the ten input languages. The testing in each round consisted of 23 memorization and 13 generalization examples. This makes a total of 2.3M memorization and 1.3M generalization test results subject to statistical analyses. Significance was tested using linear mixed-effects models, as implemented in the Python package statsmodels [100], for production similarity (LME 1), generalization score (LME 3), generalization convergence (LME 4), as well as production similarity to humans in memorization (LME 2) and generalization (LME 5). We use the structure score and the logarithmized round number in all measures as a fixed effect. The number of rounds was logarithmized following scaling laws of neural language models [60]. Both the structure score and the logarithmized round number were centered and scaled We consider two random effects: the random seed for initialization (which also determines the input language) and the specific scene. For LME 5, scaling the log-transformed round number to unit variance hindered convergence, so the log rounds were only centered. The full results of the statistical models are provided in Tab. 4 in the SI, with partial regression plots shown in Figs. 13–17. In Tab. 5 in the SI, we provide an additional analysis of production similarity to ground truth at rounds 10, 40, 70, and 100.
Code Availability
The code for reproducing our experiments is available on GitHub https://github.com/lgalke/easy2deeplearn.
@@ Acknowledgments
We thank Dota Tianai Dong, Koen de Reus, Yosef Prat, Tal Simon, Willem Zuidema, Tessa Verhoef, Mitja Nikolaus, Marieke Woensdregt, and Adam Kohan for their comments and discussions. We thank Shinje Wu and Marianne de Heer Kloots for sharing their data.
Appendix A Supplementary Information
Details of the input languages
Table 3.
@@ Appendix A Supplementary Information
Extended Results
Production Similarity to Ground Truth during Memorization
Figure 6
@@
Production Similarity to Humans during Memorization
Figure 7
@@
Production Similarity to Humans during Generalization
Figure 10
@@
Binary Accuracy with respect to Ground Truth during Memorization
Figure 11
Guessing Accuracy
Figure 12
Statistical Analyses of the Results
LME 1: Production Similarity during Memorization
LME 2: Production Similarity to Humans during Memorization
LME 3: Systematicity during Generalization
LME 4: Convergence Score during Generalization
LME 5: Production Similarity to Humans during Generalization
LME 6: Production Similarity to Ground Truth at Specific Rounds
@@
Table 4: Linear Mixed-Effects Regression Results. LME 1: Production similarity to ground truth during memorization. LME 2: Production similarity to Humans during Generalization. LME 3: Generalization Score. All tests are two-sided.
@@
Table 5: Linear Mixed-Effects Regression Results: Production Similarity to Humans during Memorization at a Fixed Round Number. All tests are two-sided.
@@
Sensitivity to Hidden Layer Size
@@
Sensitivity to the Scaling Factor for the Contrastive Loss Term
@@
Structure effect does not depend on specific hyperparameter choices
@@
Example Data
We provide example data from the memorization test in Table 6 and from the generalization test in Table 2. The sample is stratified with respect to the by-producer average production similarity to human participants. For each percentile out of 0, 25, 50, 75, and 100, we randomly sample 8 items. All samples are taken at the end of training (after epoch 100).
Table 6: Example Data from the Memorization Test.
Sim. to Humans Producer Lang. Shape Angle True Label Participant Label RNN Label
0.247826 4082 S1 4 225 wufk moof wufk
0.247826 4082 S1 2 330 weft puif weft
0.247826 4082 S1 1 240 smoogg huift smoogg
0.247826 4082 S1 4 300 suith keft suith
0.247826 4082 S1 3 135 spuut weft spuut
0.247826 4082 S1 1 60 koof puif koof
0.247826 4082 S1 4 120 wafk sgeft wafk
0.247826 4082 S1 3 270 huipt koom huipt
0.700932 1093 B2 4 270 gnt gnt gnt
0.700932 1093 B2 4 135 gnts gntsi gnts
0.700932 1093 B2 2 330 gmps gmps gmps
0.700932 1093 B2 2 120 wangsus wangsus wangsu
0.700932 1093 B2 2 45 gempt wangsus gempt
0.700932 1093 B2 4 330 gntsi gnit gntsi
0.700932 1093 B2 2 300 wangsu gmps wangsu
0.700932 1093 B2 1 30 sket sket sket
0.883282 3090 S5 2 180 nif-a nif-a nif-as
0.883282 3090 S5 1 90 wef-t wef-t wef-t
0.883282 3090 S5 1 150 wef-aat wef-att wef-aat
0.883282 3090 S5 1 240 wef-ssa wef-aas wef-ssa
0.883282 3090 S5 4 210 smut-aas smut-aas smut-aas
0.883282 3090 S5 3 180 pti-a pti-a pti-a
0.883282 3090 S5 3 315 pti-kks pti-ssk pti-kks
0.883282 3090 S5 2 225 nif-as nif-as nif-as
0.956522 1020 B1 2 330 wak-ta wak-ta wak-ta
0.956522 1020 B1 4 225 gtgt gtgt gtgt
0.956522 1020 B1 4 120 ftft ftft ftft
0.956522 1020 B1 2 150 hehi hehi hehi
0.956522 1020 B1 1 210 ha-ia ha-ia ha-ia
0.956522 1020 B1 3 45 fiti fiti fiti
0.956522 1020 B1 4 60 kite fik kite
0.956522 1020 B1 4 360 pepepe pepepe pepepe
1.000000 1008 B4 4 225 fak-huif fak-huif fak-huif
1.000000 1008 B4 3 45 muif-a muif-a muif-a
1.000000 1008 B4 4 120 fak-e fak-e fak-e
1.000000 1008 B4 1 315 fas-pok fas-pok fas-pok
1.000000 1008 B4 3 135 muif-e muif-e muif-e
1.000000 1008 B4 2 135 pok-e pok-e pok-e
1.000000 1008 B4 4 300 fak-pok fak-pok fak-pok
1.000000 1008 B4 2 210 pok-huif pok-huif pok-huif
Table 7: Example Data from the Generalization Test.
Sim. to Humans Producer Lang. Shape Angle Participant Label RNN Label
0.092308 1048 B1 2 360 kokoke seefe
0.092308 1048 B1 1 225 po-ti ha-ia
0.092308 1048 B1 3 90 ghio mimi
0.092308 1048 B1 4 240 khio gtgt
0.092308 1048 B1 3 150 ptpt mimi
0.092308 1048 B1 2 300 ko-toe wak-ta
0.092308 1048 B1 4 210 ka-ia gtgt
0.092308 1048 B1 2 225 haia pooti
0.443223 8064 B2 3 150 wangsi wangsuu
0.443223 8064 B2 4 225 gntsoe gntuu
0.443223 8064 B2 1 135 sketsi gesh
0.443223 8064 B2 4 360 gnt skek
0.443223 8064 B2 4 60 gmpsi skek
0.443223 8064 B2 2 270 wng wangsuu
0.443223 8064 B2 2 360 wang gempt
0.443223 8064 B2 3 225 wangsuus wangsoe
0.590884 1083 S3 4 60 fuottee fuoto-o-o-o
0.590884 1083 S3 4 150 fuottoo fuottii
0.590884 1083 S3 1 30 fewo-o-o-o fewen
0.590884 1083 S3 3 60 powi powu-u-u
0.590884 1083 S3 1 225 fewo-o-o-o fewo-o-o
0.590884 1083 S3 3 225 powee powwoo
0.590884 1083 S3 2 360 asken asko-o-o
0.590884 1083 S3 4 330 fuottoa fuotio
0.772497 2058 B4 2 360 pok-i pok
0.772497 2058 B4 4 330 fak-pok fas-i
0.772497 2058 B4 4 60 fak-a fak-e
0.772497 2058 B4 1 30 fas-a fas-a
0.772497 2058 B4 2 300 pok pok-u
0.772497 2058 B4 4 90 fak-u fak-e
0.772497 2058 B4 4 150 fak-w-w-e fak-e
0.772497 2058 B4 3 225 muif-huif muif-huif
0.923443 5016 B4 4 90 fak-iii fak-e
0.923443 5016 B4 1 120 fas-e fas-e
0.923443 5016 B4 2 300 pok pok
0.923443 5016 B4 3 60 muif-a muif-a
0.923443 5016 B4 3 360 muif-i muif-a
0.923443 5016 B4 4 150 fak-e fak-e
0.923443 5016 B4 3 225 muif-huif muif-huif
0.923443 5016 B4 4 150 fak-e fak-e